# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

### v4

@@ 1 Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2022), PaLM (Chowdhery et al., 2023), Llama (Grattafiori et al., 2024) and DeepSeek (Bi et al., 2024) have revolutionized the field of generative artificial intelligence by enabling large-scale content generation across various fields including journalism, education, and creative writing (Demszky et al., 2023; Milano et al., 2023; Doshi & Hauser, 2024). However, their ability to produce highly human-like text poses serious risks, such as the spread of misinformation, academic dishonesty, and the erosion of trust in written communication (Ahmed et al., 2021; Lee et al., 2023; Christian, 2023). Consequently, accurately distinguishing between human- and LLM-generated text has emerged as a critical area of research.
@@ 1.1 Related works
In addition to these two types of methods, Abburi et al. (2023) propose a hybrid approach that uses the outputs of fine-tuned LLMs as input features for classical ML-based classification. Further efforts have focused on handling adversarial attacks (Crothers et al., 2022; Krishna et al., 2023; Koike et al., 2024; Sadasivan et al., 2025), short texts (Tian et al., 2024), out-of-distribution texts Guo et al. (2024b), biases against non-native English writers (Liang et al., 2023), as well as the downstream applications of these methods in domains such as education, social media and medicine (Herbold et al., 2023; Kumarage et al., 2023; Liao et al., 2023).
@@ 2 Preliminaries
We first define the white-box and black-box settings as well as our objective. We next review two baseline methods, DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024), as they are closely related to our proposal. Finally, we introduce the martingale central limit theorem (see e.g., Hall & Heyde, 2014), which serves as the theoretical basis for our threshold selection.
@@ 3 AdaDetectGPT
  • Theorem 2 provides an upper bound on the difference between our classifier’s expected FNR and the target level $\alpha$. This excess FNR depends on two factors: (i) the number of tokens $L$, and (ii) the variance ratio $V_{L}$. As $L$ diverges to infinity and $V_{L}$ stabilizes to $1$, the FNR of our classifier approaches the target level. We remark that assumptions similar to the bounded denominator condition in Theorem 2 are commonly imposed (see e.g., Bolthausen, 1982; Hall & Heyde, 2014).
@@ 5 Conclusion
In our implementation, we opted to learn the witness function via a B-spline basis due to the straightforward nature of the optimization (finding the optimal witness function boils down to solving a system of linear equations) and its favorable theoretical properties (the estimation error can attain Stone’s optimal convergence rate, Stone, 1982, see Appendix E.6 for more details). The number of basis functions can be selected in a data driven way via Lepski’s method (Lepski & Spokoiny, 1997).
@@ Appendix A Additional related works: Watermarking-based detection
Watermarking embeds subtle signals into LLM-generated text to distinguish it from human-written text (see Ji et al., 2025, Section 4.2, for a recent review of LLM watermarking). Aaronson & Kirchner (2023) propose a watermarking technique based on Gumbel sampling. Follow-up works have focused on preserving text quality during watermarking (Christ et al., 2024; Dathathri et al., 2024; Giboulot & Furon, 2024; Liu & Bu, 2024; Wouters, 2024; Wu et al., 2024), enhancing watermark detection (Dathathri et al., 2024; Huo et al., 2024; Cai et al., 2025) and maintaining robustness against adversarial edits (Golowich & Moitra, 2024).
Our work is related to a line of research that frames watermark detection as a statistical hypothesis testing problem (see e.g., Kirchenbauer et al., 2023; Hu et al., 2024; Kuditipudi et al., 2024; Zhao et al., 2024; Li et al., 2025a, b; Chen et al., 2025b). Under this framework, rejection of the null hypothesis (that no watermark is present) provides statistical evidence that the text was likely generated by an LLM.
@@ Appendix C Experiment details
$\widehat{\beta}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\widehat{\alpha}$,
@@
GPT-J (Wang & Komatsuzaki, 2021) EleutherAI/gpt-j-6B 6B
@@ Appendix D Technical assumptions
Conditions similar to Assumption 1 are commonly assumed; see, for instance, Audibert & Tsybakov (2007); Qian & Murphy (2011); Luedtke & Van Der Laan (2016); Shi et al. (2020a, b, 2022). Assumption 2 is mild, since the constant $\gamma$ is allowed to be arbitrarily large. Assumption 3 basically requires the conditional variance of logits be asymptotically equivalent for human-authored and machine-generated passages. This assumption is not overly restrictive, as the variance discrepancy between the two types of passages is relatively small in our dataset (see Table S5, where the ratio f the variances are very closed to 1). Assumption 4 is commonly assumed in martingale central limit theorem literature, see e.g. Bolthausen (1982); Hall & Heyde (2014). Our empirical results further support the validity of Assumption 4: the sample mean of the ratio remains nearly constant as a function of $L$, while its variance (in parentheses) approaches zero as $L\to\infty$ across all three datasets (see Table S5), suggesting that this condition is practical. Furthermore, two commonly employed hypothesis tests — the Kolmogorov–Smirnov (KS) test and the Shapiro–Wilk (SW) test — are conducted to evaluate whether the proposed statistic follows a normal distribution. As shown in Table S6, almost all $p$-values exceed 0.1 (most by a large margin), indicating that our test statistic passes normality tests in most cases. These test results also provide strong empirical support for the validity of MCLT regularity conditions.
@@ E.2 Preparatory results
See Corollary 3.1 in Hall & Heyde (2014) and Theorem 2 in Bolthausen (1982). ∎
@@ E.6 Smooth witness functions
An important quality of spline estimators, which in part motivated our choice of estimator for the witness function, is their ability to learn smooth regression functions at optimal rates. Following the proof of Theorem 2.1 in Chen & Christensen (2015) one can show that when the optimal witness (that is, the function minimizing the functional (21) among all functions with bounded $\ell_{2}$ norm) is $\beta$-Hölder smooth, the estimated witness function attains the rate
@@
where (41a) follows from the arguments leading up to (34) and (41b) holds on the event (35). Using the mean value theorem and Assumption 3 one can show that up to constants $\hat{\Delta}_{n}$ is smaller than $\sup_{z}\left|\widehat{w}(z)-w^{*}(z)\right|$, therefore by (40) the first term in (41b) will be of the order $O(d^{-\beta}+\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Examining the proof of Theorem 2.1 in Chen & Christensen (2015) it can be seen that for any $\kappa_{3}>0$ on may choose $\kappa_{1}$ so that $\mathbb{P}(\neg A(\kappa_{1}))<n^{-\kappa_{3}}$. Therefore, we can choose $\kappa_{1}$ appropriately so that the second term in (41b) will be of the order $o(\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Finally letting $d=\Theta((n/\log(n))^{\frac{1}{2\beta+1}})$ yields the desired result.
As mentioned in the main text, when the optimal witness function is believed to be smooth but the order of the smoothness is not known the number of spline bases can be chosen in a data driven way via Lepski’s method (Lepski & Spokoiny, 1997).

### v5

@@ 1 Introduction
Large language models (LLMs) such as ChatGPT (OpenAI, 2022), PaLM (Chowdhery et al., 2023), Llama (Grattafiori et al., 2024) and DeepSeek (Bi et al., 2024) have revolutionized the field of generative artificial intelligence by enabling large-scale content generation across various fields including journalism, education, and creative writing (Demszky et al., 2023; Milano et al., 2023; Doshi and Hauser, 2024). However, their ability to produce highly human-like text poses serious risks, such as the spread of misinformation, academic dishonesty, and the erosion of trust in written communication (Ahmed et al., 2021; Lee et al., 2023; Christian, 2023). Consequently, accurately distinguishing between human- and LLM-generated text has emerged as a critical area of research.
@@ 1.1 Related works
In addition to these two types of methods, Abburi et al. (2023) propose a hybrid approach that uses the outputs of fine-tuned LLMs as input features for classical ML-based classification. Further efforts have focused on handling adversarial attacks (Crothers et al., 2022; Krishna et al., 2023; Koike et al., 2024; Sadasivan et al., 2025), short texts (Tian et al., 2024), out-of-distribution texts (Guo et al., 2024b), unobserved prompts (Zhou et al., 2026a), biases against non-native English writers (Liang et al., 2023), accommodating statistical inference (Zhou et al., 2026b), as well as the downstream applications of these methods in domains such as education, social media and medicine (Herbold et al., 2023; Kumarage et al., 2023; Liao et al., 2023).
@@ 2 Preliminaries
We first define the white-box and black-box settings as well as our objective. We next review two baseline methods, DetectGPT (Mitchell et al., 2023) and Fast-DetectGPT (Bao et al., 2024), as they are closely related to our proposal. Finally, we introduce the martingale central limit theorem (see e.g., Hall and Heyde, 2014), which serves as the theoretical basis for our threshold selection.
@@ 3 AdaDetectGPT
  • Theorem 2 provides an upper bound on the difference between our classifier’s expected FNR and the target level $\alpha$. This excess FNR depends on two factors: (i) the number of tokens $L$, and (ii) the variance ratio $V_{L}$. As $L$ diverges to infinity and $V_{L}$ stabilizes to $1$, the FNR of our classifier approaches the target level. We remark that assumptions similar to the bounded denominator condition in Theorem 2 are commonly imposed (see e.g., Bolthausen, 1982; Hall and Heyde, 2014).
@@ 5 Conclusion
In our implementation, we opted to learn the witness function via a B-spline basis due to the straightforward nature of the optimization (finding the optimal witness function boils down to solving a system of linear equations) and its favorable theoretical properties (the estimation error can attain Stone’s optimal convergence rate, Stone, 1982, see Appendix E.6 for more details). The number of basis functions can be selected in a data driven way via Lepski’s method (Lepski and Spokoiny, 1997).
@@ Appendix A Additional related works: Watermarking-based detection
Watermarking embeds subtle signals into LLM-generated text to distinguish it from human-written text (see Ji et al., 2025, Section 4.2, for a recent review of LLM watermarking). Aaronson and Kirchner (2023) propose a watermarking technique based on Gumbel sampling. Follow-up works have focused on preserving text quality during watermarking (Christ et al., 2024; Dathathri et al., 2024; Giboulot and Furon, 2024; Liu and Bu, 2024; Wouters, 2024; Wu et al., 2024), enhancing watermark detection (Dathathri et al., 2024; Huo et al., 2024; Cai et al., 2025) and maintaining robustness against adversarial edits (Golowich and Moitra, 2024).
Our work is related to a line of research that frames watermark detection as a statistical hypothesis testing problem (see e.g., Kirchenbauer et al., 2023; Hu et al., 2024; Kuditipudi et al., 2024; Zhao et al., 2024; Li et al., 2025b, a; Chen et al., 2025b). Under this framework, rejection of the null hypothesis (that no watermark is present) provides statistical evidence that the text was likely generated by an LLM.
@@ Appendix C Experiment details
$\widehat{\beta}=(\widehat{\Sigma}^{(h)}+\widehat{\Sigma}^{(m)})^{-1/2}\widehat{\alpha}$.
@@
GPT-J (Wang and Komatsuzaki, 2021) EleutherAI/gpt-j-6B 6B
@@ Appendix D Technical assumptions
Conditions similar to Assumption 1 are commonly assumed; see, for instance, Audibert and Tsybakov (2007); Qian and Murphy (2011); Luedtke and Van Der Laan (2016); Shi et al. (2020b, a, 2022). Assumption 2 is mild, since the constant $\gamma$ is allowed to be arbitrarily large. Assumption 3 basically requires the conditional variance of logits be asymptotically equivalent for human-authored and machine-generated passages. This assumption is not overly restrictive, as the variance discrepancy between the two types of passages is relatively small in our dataset (see Table S5, where the ratio f the variances are very closed to 1). Assumption 4 is commonly assumed in martingale central limit theorem literature, see e.g. Bolthausen (1982); Hall and Heyde (2014). Our empirical results further support the validity of Assumption 4: the sample mean of the ratio remains nearly constant as a function of $L$, while its variance (in parentheses) approaches zero as $L\to\infty$ across all three datasets (see Table S5), suggesting that this condition is practical. Furthermore, two commonly employed hypothesis tests — the Kolmogorov–Smirnov (KS) test and the Shapiro–Wilk (SW) test — are conducted to evaluate whether the proposed statistic follows a normal distribution. As shown in Table S6, almost all $p$-values exceed 0.1 (most by a large margin), indicating that our test statistic passes normality tests in most cases. These test results also provide strong empirical support for the validity of MCLT regularity conditions.
@@ E.2 Preparatory results
See Corollary 3.1 in Hall and Heyde (2014) and Theorem 2 in Bolthausen (1982). ∎
@@ E.6 Smooth witness functions
An important quality of spline estimators, which in part motivated our choice of estimator for the witness function, is their ability to learn smooth regression functions at optimal rates. Following the proof of Theorem 2.1 in Chen and Christensen (2015) one can show that when the optimal witness (that is, the function minimizing the functional (21) among all functions with bounded $\ell_{2}$ norm) is $\beta$-Hölder smooth, the estimated witness function attains the rate
@@
where (41a) follows from the arguments leading up to (34) and (41b) holds on the event (35). Using the mean value theorem and Assumption 3 one can show that up to constants $\hat{\Delta}_{n}$ is smaller than $\sup_{z}\left|\widehat{w}(z)-w^{*}(z)\right|$, therefore by (40) the first term in (41b) will be of the order $O(d^{-\beta}+\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Examining the proof of Theorem 2.1 in Chen and Christensen (2015) it can be seen that for any $\kappa_{3}>0$ on may choose $\kappa_{1}$ so that $\mathbb{P}(\neg A(\kappa_{1}))<n^{-\kappa_{3}}$. Therefore, we can choose $\kappa_{1}$ appropriately so that the second term in (41b) will be of the order $o(\left(\log(n)/n\right)^{\frac{\beta}{2\beta+1}})$. Finally letting $d=\Theta((n/\log(n))^{\frac{1}{2\beta+1}})$ yields the desired result.
As mentioned in the main text, when the optimal witness function is believed to be smooth but the order of the smoothness is not known the number of spline bases can be chosen in a data driven way via Lepski’s method (Lepski and Spokoiny, 1997).